# Characterizing Pattern Matching and Its Limits on Compositional Task Structures

### v2

@@ 1 Introduction
Despite the remarkable performance of Large Language Models (LLMs) (brown2020language; touvron2023llama), compositional generalization studies suggest that the core mechanism of generalization might be “pattern matching”, i.e., models learning local statistical regularities between input fragments and outputs in some cases (loula2018rearranging; johnson2017clevr; berglund2023reversal; wang2024grokked; mirzadeh2024gsm; keysers2019measuring; csordas2022ctl++). However, behavioral studies commonly employ task setups that allow multiple generalization sources (e.g., algebraic invariances, structural repetition), discussing pattern matching without a precise definition and diagnosing it post-hoc rather than characterizing it predictively. As a result, it remains unclear which behaviors should count as pattern matching and which should not, obscuring a constructive and testable account of its boundary.
@@
To make this notion precise, we (1) introduce a model-agnostic, data-centric formalism for pattern matching and (2) systematically study how modern architectures like decoder-only Transformers (vaswani2017attention) and Mamba (gu2023mamba) perform generalization through pattern matching. Specifically, we first propose a model-agnostic and data-centric definition of pattern matching by formalizing the substitution of input patterns observed to result in identical outputs in shared contexts as functional-equivalence (Sec.˜3; henceforth, we use pattern matching as equivalent to functional-equivalence-based generalization). This induces a coverage boundary: if learning relies only on such evidence, reliable prediction is expected only for test inputs reachable by these substitutions.
@@
Moreover, to isolate and study pattern-matching behaviors, we use controlled setups that deliberately remove other generalization sources and make functional equivalence the primary available mechanism. With this setting, our formalism yields predictive and quantitative insights about the limitations of pattern matching that, to our knowledge, have not been well characterized in prior works:
@@
  • Generalization success is well predicted by the number of supporting contexts that witness the relevant functional equivalence. Mechanistically, Transformers implement functional equivalence via clustered intermediate representations at specific layers/positions, with clustering strength aligning with evidence strength (Sec.˜5).
@@
  • We prove a tight sample complexity bound of pattern matching on a two-hop structure by identifying the exponent of the data scaling law (in terms of the token set size) which is necessary and sufficient for perfect in-domain generalization (Theorem˜6.1). The measured power-law exponent agrees with our theoretical bounds and remains stable under roughly $20\times$ parameter increase (from 68M to 1.5B) for GPT-2 (radford2019language), and also holds for Mamba (gu2023mamba) architecture (Sec.˜6).
@@
  • When the same variable influences the output along multiple computational paths, models fail to form unified intermediate state representations. Analysis reveals that they instead develop context-dependent state representations, impairing both generalization and interpretability (Sec.˜7).
@@
  • Chain-of-Thought (CoT) supervision (wei2022chain) reduces data requirements yet does not resolve path ambiguity without seeing nearly exhaustive in-domain combinations (Sec.˜8).
@@ 2 Related Work
Pattern matching behaviors of LLMs on compositional tasks.   It is well perceived that pattern matching alone is inadequate for systematic generalization (fodor1988connectionism), and modern LLMs display generalization abilities that seem to be far beyond what pattern matching alone can do, as measured by their remarkable performance on complex benchmarks (achiam2023gpt). However, a growing body of work has consistently reported that LLMs still fall short on benchmarks designed to test compositionality (hupkes2020compositionality), including mathematical reasoning (mirzadeh2024gsm), multi-hop reasoning (yang2024large; wang2024grokked), and more (lake2018generalization; kim2020cogs; csordas2022ctl++; dziri2024faith). This gap between their capabilities and pattern-matching behaviors on compositional tasks calls for a principled framework to define what pattern matching is and to what extent a model’s behavior can be attributed to pattern matching, but it is mostly discussed with behavioral studies under the context of a specifically designed benchmark. Our work addresses this gap by formally defining pattern matching, and we systematically analyze models’ behaviors with controlled tasks that are designed to isolate pattern-matching regimes grounded in our framework.
Mechanistic interpretability.   Mechanistic interpretability studies aim to understand how submechanisms implement models’ behaviors (elhage2021mathematical; olsson2022context; nanda2023progress; elhage2022toy). Recent work analyzes how Transformer components are causally related to certain behaviors (meng2022locating; hanna2023does; goldowsky2023localizing). In particular, it is reported that in-domain compositional generalization can emerge through grokking, with identifiable intermediate state representations inside Transformers (wang2024grokked). Our framework complements these works by providing mechanistic insights about pattern matching. Our findings also explain why standard interpretability techniques like logit lens (nostalgebraist2020logitlens; belrose2023eliciting) may fail to identify state representations in models trained on tasks with path ambiguities.
@@ 3 Formalizing Pattern Matching with Functional Equivalence
Imagine a learner observing data determined by $f:{\mathcal{X}}^{3}\to{\mathcal{X}}$. The input ${\bm{x}}=(x_{1},x_{2},x_{3})\in{\mathcal{X}}^{3}$ is a sequence of three discrete tokens and the output is a single token, where each token is chosen from a finite set $\mathcal{X}$}. [^1] For brevity, we use a shared token set $\mathcal{X}$} in the main text. A more general notion using position-specific domains (e.g., ${\mathcal{X}}_{i}$’) can be used; see Sec. F.1. Suppose (unknown to the learner) that $f$ factorizes as the composition of two primitive functions, $f({\bm{x}})=f_{2}(f_{1}(x_{1},x_{2}),x_{3})$, where $f_{1}:{\mathcal{X}}^{2}\to{\mathcal{X}}$ and $f_{2}:{\mathcal{X}}^{2}\to{\mathcal{X}}$, as illustrated in Fig.˜2(a). How can the learner generalize by only seeing the input-output patterns?
@@
Fix a nonempty proper subset $I$ of indices in $\ell$]. Consider any set $S\subset{\mathcal{X}}^{\ell}$ of input sequences. [^2] The set $S$ can be any subset of the whole domain, e.g., $\mathcal{X}$$^{\ell}$ itself, the train dataset $D$, or whatever else. Given a pair of subsequences ${\bm{a}},{\bm{a}}^{\prime}\in{\mathcal{X}}^{\left|I\right|}$, we say a pair of inputs $\left\{{\bm{x}},{\bm{x}}^{\prime}\right\}$ to be an $I$-co-occurrence of $\bm{a}$} and $\bm{a}$$^{\prime}$ in $S$ if it satisfies $\left\{{\bm{x}},{\bm{x}}^{\prime}\right\}\subset S$ and ${\bm{x}}_{I}={\bm{a}}$, ${\bm{x}}^{\prime}_{I}={\bm{a}}^{\prime}$, and ${\bm{x}}_{[\ell]\setminus I}={\bm{x}}^{\prime}_{[\ell]\setminus I}$. Also, the subsequences $\bm{a}$} and $\bm{a}$$^{\prime}$ are said to be functionally $k$-equivalent at $I$ in $S$ and denoted by ${\bm{a}}\equiv^{I}_{S}{\bm{a}}^{\prime}$, if it satisfies:
@@
Next, we ask: which inputs are reachable from observed data utilizing functional equivalence? To formalize this, we define substitution graph: Let ${\mathcal{G}}^{(D,k)}=(V,E)$ be an undirected graph with a vertex set $V={\mathcal{X}}^{\ell}$ of all possible inputs. Two vertices ${\bm{x}},{\bm{x}}^{\prime}\in V$ are connected with an edge in $E$ if and only if there exists an index set $I\subset[\ell]$ such that $\left\{{\bm{x}},{\bm{x}}^{\prime}\right\}$ is an $I$-co-occurrence (in $V$) of a pair of functionally $k$-equivalent sequences at $I$ in $D$. This process is illustrated on the right side of Fig.˜1, as a special case where $k=1$. With this substitution graph ${\mathcal{G}}^{(D,k)}$, we formally define the $k$-coverage as a set of inputs which are connected [^3] For an undirected graph $\mathcal{G}$}, two vertices $u,v$ are connected if $\mathcal{G}$} contains a path between $u$ and $v$. to at least one observed input as follows:
@@
Note that the notion of coverage is a stricter condition of the canonical definition of in-domain (ID), which is obtained by random train/test split (wang2024grokked) or taking combinations of observed internal computations (dziri2024faith). In Sec.˜5, we demonstrate that learners may not necessarily generalize on data that are classified as ID in a canonical sense, but coverage can precisely explain when and why this occurs. We also emphasize that coverage is a property of a dataset and is independent of model architectures and learning algorithms, and we demonstrate that the predictions made by our framework are invariant across model architecture and scale in Sec.˜6 and 7. Finally, $k$-coverage can be algorithmically determined for any fixed-length discrete sequence tasks (Alg. 1), which we use for the analyses in the following sections.
@@ 4 Experimental Setup
Dataset construction.   We construct four synthetic tasks with different structures: 2-Hop, Parallel 2-Hop, 3-Hop, and Non-tree (Fig.˜2). To isolate functional-equivalence-based generalizations, we create random mappings from a product space of token sets to control the generalization sources not attributable to compositional structures (i.e., commutativity). We explain the dataset construction process using 2-Hop task (Fig.˜2(a)), $(x_{1},x_{2},x_{3})\mapsto t$ with $t=f_{2}(f_{1}(x_{1},x_{2}),x_{3})$, as an example. We construct training datasets by defining a token set with size $\left|{\mathcal{X}}\right|$, and creating two random maps for the primitive functions $f_{1}:{\mathcal{X}}^{2}\to{\mathcal{X}}$ and $f_{2}:{\mathcal{X}}^{2}\to{\mathcal{X}}$. We mark a fraction $p_{\mathrm{seen}}=0.7$ of each function’s domain as ‘seen’, gather all possible combinations where both functions are applied to inputs from their seen domains, and uniformly sample $N$ examples to form a training dataset. See Sec.˜B.1 for more details of the dataset construction process.
Training & evaluation.   Following wang2022interpretability, we train randomly initialized GPT-2 (radford2019language) models with 8 layers, 12 heads, and 768 dimensions as a base model (see Sec.˜B.2 for details). We construct two evaluation sets, each with 2,000 instances: (1) ID Test Set : all primitive function applications (e.g., $f_{1}(x_{1},x_{2})$ and $f_{2}(b,x_{3})$ in 2-Hop task) are observed during training, but their specific combination was unseen. (2) Out-of-coverage (canonical OOD) Test Set: at least one primitive function application is never observed during training, which is used as a control group.
@@ 5.1 Evidence Strength is Tightly Aligned to Pattern-Matching Success
We first analyze the correlation between $k$-coverage and ID generalization performance of the GPT-2 model. To this end, we implement and release a task-agnostic coverage determination algorithm (see App.˜C) that can be applied to diverse compositional structures. Then, we analyze what fraction of ID test data of 2-Hop task with $\left|{\mathcal{X}}\right|=50$ lies inside $k$-coverage, depending on $k$ and dataset size $N$. Fig.˜3 (Left) shows that at N=5N=5k, every ID test example is already covered with minimal evidence ($k=1$). Hence, in an ideal scenario where a single witness of functional equivalence suffices, training with the dataset as small as N=5N=5k will lead to perfect ID generalization.
@@
These results yield two important insights. First, successful ID generalization in practice requires a robust coverage so the model can confidently identify and utilize functional equivalence relationships. The parameter $k$ effectively quantifies this evidence strength, directly impacting generalization speed and reliability. Second, while our experiments use uniformly sampled datasets, the results can explain why models struggle with generalizing long-tail distributions in imbalanced real-world data (mallen2022not; kandpal2023large; chang2024large). Despite technically being in-distribution, rare combinations naturally receive limited evidence of functional equivalence (low $k$), effectively and practically placing them outside the coverage. We believe that our insights will motivate future research on targeted data augmentation strategies to maximize $k$-coverage.
@@ 5.2 Latent Representation Clusters Drive Pattern Matching on Coverage
To quantify this representation clustering phenomenon, we develop a metric that captures how distinctly the model separates functionally equivalent fragments from others. Specifically, we measure the difference between the average pairwise cosine similarity of latent vectors that share the same intermediate state $b=f_{1}(x_{1},x_{2})$ ($\overline{\cos}_{\text{intra}}$), and those that do not ($\overline{\cos}_{\text{inter}}$), for each position and layer of the model. We term this difference the intra–inter cosine gap $\text{IICG}=\overline{\cos}_{\text{intra}}-\overline{\cos}_{\text{inter}}$, where higher values indicate stronger within-group clustering relative to between-group separation. Fig.˜4 (Left) reveals a positive correlation: higher $k$-cutoff values yield higher IICG scores at certain positions, indicating that stronger functional equivalence evidence leads to more coherent internal representations. In contrast, out-of-coverage ($k=0$) examples exhibit no clustering pattern, as they lack evidence of functional equivalence in the training data. The PCA visualization at position $x_{2}$ and layer 3 (Right) shows this trend visually. We verify that the representation clusters play a causal role in pattern matching with causal tracing (goldowsky2023localizing; hanna2023does), a widely used technique to identify Transformer circuits (Fig.˜8).
Our findings extend the previous insights from mechanistic interpretability literature (wang2024grokked) in several ways. First, we demonstrate that unified circuit formation is driven by functional equivalence evidence in the training data, not by explicit exposure to intermediate computation steps. Moreover, we find that these clustered representations are not necessarily aligned with vocabulary embeddings, implying that standard interpretability methods like logit lens nostalgebraist2020logitlens may fail to detect these functional equivalence representations despite their presence (see App.˜E).

### v3

# arXiv:2505.20278v3[cs.LG] 02 Mar 2026

@@ 1 Introduction
Despite the remarkable performance of large language models (LLMs) (Brown et al., 2020; Touvron et al., 2023), compositional generalization studies suggest that the core mechanism of generalization might be “pattern matching,” i.e., models learning local statistical regularities between input fragments and outputs in some cases (Loula et al., 2018; Johnson et al., 2017; Berglund et al., 2024; Wang et al., 2024a; Mirzadeh et al., 2025; Keysers et al., 2020; Csordás et al., 2022). However, behavioral studies commonly employ task setups that allow multiple generalization sources (e.g., algebraic invariances, structural repetition), discussing pattern matching without a precise definition and diagnosing it post-hoc rather than characterizing it predictively. As a consequence, it remains unclear which behaviors should be counted as pattern matching and which should not, thereby obscuring a constructive and testable account of its boundary.
@@
To make this notion precise, we (1) introduce a data-centered formalism for pattern matching and (2) systematically study how modern architectures like decoder-only Transformers (Vaswani et al., 2017; Radford et al., 2018; 2019) and Mamba (Gu and Dao, 2024) perform generalization through pattern matching. Specifically, we first propose a model-agnostic and data-centric definition of pattern matching by formalizing the substitution of input patterns observed to result in identical outputs in shared contexts as functional-equivalence (Def.˜3.1; henceforth, we use pattern matching as equivalent to functional-equivalence-based generalization). This induces a coverage boundary (Def.˜3.2): if learning relies only on such evidence, there is no hope to expect any reliable predictions for the test inputs unreachable by these substitutions.
@@
Moreover, to isolate and study pattern-matching behaviors, we use controlled setups that deliberately remove other sources of generalization and make functional equivalence the primary available mechanism. With these setups, our formalism offers predictive and quantitative insights about the limitations of pattern matching. To the best of our knowledge, these have not been well characterized yet in prior works:
@@
  • The success in generalization is well predicted by the number of supporting contexts that witness the relevant functional equivalence. Mechanistically, Transformers seem to implement the functional equivalence by clustering intermediate representations at specific layers/positions. The strength of the clustering aligns with the evidence strength (Sec.˜5).
@@
  • We prove a tight sample complexity bound of pattern matching on a two-hop structure by identifying the exponent of the data scaling law (in terms of the token set size) which is necessary and sufficient to achieve perfect in-domain generalization (Theorem˜6.1). We find that the measured power-law exponent agrees with our theoretical bounds, remains stable under roughly $20\times$ parameter increase (from 68M to 1.5B) for GPT-2 (Radford et al., 2019), and also holds for Mamba (Gu and Dao, 2024) architecture (Sec.˜6).
@@
  • When the same variable influences the output along multiple computational paths, models fail to form unified intermediate state representations. Our analysis reveals that they instead develop context-dependent state representations, impairing both generalization and interpretability (Sec.˜7).
@@
  • Chain-of-Thought (CoT) supervision (Wei et al., 2022) does not completely resolve the path ambiguity without observing nearly exhaustive in-domain combinations, although it reduces the data requirements to some extent (Sec.˜8).
@@ 2 Related Work
Pattern matching behaviors of LLMs on compositional tasks.   It is well perceived that pattern matching alone is inadequate for systematic generalization (Fodor and Pylyshyn, 1988), and modern LLMs display generalization abilities that seem to be far beyond what pattern matching alone can do, as measured by their remarkable performance on complex benchmarks (OpenAI, 2023). However, a growing body of work has consistently reported that LLMs still fall short on benchmarks designed to test compositionality (Hupkes et al., 2020), including mathematical reasoning (Mirzadeh et al., 2025), multi-hop reasoning (Yang et al., 2024; Wang et al., 2024a), and more (Lake and Baroni, 2018; Kim and Linzen, 2020; Csordás et al., 2022; Dziri et al., 2023). This gap between their capabilities and pattern-matching behaviors on compositional tasks calls for a principled framework to define what pattern matching is and to what extent a model’s behavior can be attributed to pattern matching, but it is mostly discussed with behavioral studies under the context of a specifically designed benchmark. Our work addresses this gap by formally defining pattern matching, and we systematically analyze models’ behaviors with controlled tasks that are designed to isolate pattern-matching regimes grounded in our framework.
Mechanistic interpretability.   Mechanistic interpretability studies aim to understand how submechanisms implement models’ behaviors (Elhage et al., 2021; Olsson et al., 2022; Nanda et al., 2023; Elhage et al., 2022). Recent work analyzes how Transformer components are causally related to certain behaviors (Meng et al., 2022; Hanna et al., 2023; Goldowsky-Dill et al., 2023). In particular, it is reported that in-domain compositional generalization can emerge through grokking, with identifiable intermediate state representations inside Transformers (Wang et al., 2024a). Our framework complements these works by providing mechanistic insights about pattern matching. Our findings also explain why standard interpretability techniques like logit lens (nostalgebraist, 2020; Belrose et al., 2023) may fail to identify state representations in models trained on tasks with path ambiguities.
@@ 3 Formalizing Pattern Matching with Functional Equivalence
Imagine a learner observing data determined by $f:{\mathcal{X}}^{3}\to{\mathcal{X}}$. The input ${\bm{x}}=(x_{1},x_{2},x_{3})\in{\mathcal{X}}^{3}$ is a sequence of three discrete tokens and the output is a single token, where each token is chosen from a finite set $\mathcal{X}$}. [^1] For brevity, we use a shared token set $\mathcal{X}$} in most parts of this paper. A more general notion using position-specific domains (e.g., ${\mathcal{X}}_{i}$’) can be used; e.g., see Sec. F.1. Suppose (unknown to the learner) that $f$ factorizes as the composition of two primitive functions, $f({\bm{x}})=f_{2}(f_{1}(x_{1},x_{2}),x_{3})$, where $f_{1}:{\mathcal{X}}^{2}\to{\mathcal{X}}$ and $f_{2}:{\mathcal{X}}^{2}\to{\mathcal{X}}$, as illustrated in Fig.˜2(a). How can the learner generalize by only seeing the input-output patterns?
@@
Fix a nonempty proper subset $I$ of indices in $\ell$]. Consider any set $S\subset{\mathcal{X}}^{\ell}$ of input sequences. [^2] The set $S$ can be any subset of the whole domain, e.g., $\mathcal{X}$$^{\ell}$ itself, the train dataset $D$, or whatever else. Given a pair of subsequences ${\bm{a}},{\bm{a}}^{\prime}\in{\mathcal{X}}^{\left|I\right|}$, we say a pair of inputs $\left\{{\bm{x}},{\bm{x}}^{\prime}\right\}$ to be an $I$-co-occurrence of $\bm{a}$} and $\bm{a}$$^{\prime}$ in $S$ if it satisfies $\left\{{\bm{x}},{\bm{x}}^{\prime}\right\}\subset S$ and ${\bm{x}}_{I}={\bm{a}}$, ${\bm{x}}^{\prime}_{I}={\bm{a}}^{\prime}$, and ${\bm{x}}_{[\ell]\setminus I}={\bm{x}}^{\prime}_{[\ell]\setminus I}$. Also, the subsequences $\bm{a}$} and $\bm{a}$$^{\prime}$ are said to be functionally $k$-equivalent at $I$ in $S$ and denoted by ${\bm{a}}\equiv^{(D,k)}_{I}{\bm{a}}^{\prime}$, if it satisfies:
@@
Next, we ask: which inputs are reachable from observed data utilizing functional equivalence? To formalize this, we define substitution graph: Let ${\mathcal{G}}^{(D,k)}=(V,E)$ be an undirected graph with a vertex set $V={\mathcal{X}}^{\ell}$ of all possible inputs. Two vertices ${\bm{x}},{\bm{x}}^{\prime}\in V$ are connected with an edge in $E$ if and only if there exists an index set $I\subset[\ell]$ such that ${\bm{x}}_{I}\equiv^{(D,k)}_{I}$ $\left\{{\bm{x}},{\bm{x}}^{\prime}\right\}$ is an $I$-co-occurrence (in $V$) of a pair of functionally $k$-equivalent sequences at $I$ in $D$. This process is illustrated on the right side of Fig.˜1, as a special case where $k=1$. With this substitution graph ${\mathcal{G}}^{(D,k)}$, we formally define the $k$-coverage as a set of inputs which are connected [^3] For an undirected graph $\mathcal{G}$}, two vertices $u,v$ are connected if $\mathcal{G}$} contains a path between $u$ and $v$. to at least one observed input as follows:
@@
Note that the notion of coverage is a stricter condition of the canonical definition of in-domain (ID), which is obtained by random train/test split (Wang et al., 2024a) or taking combinations of observed internal computations (Dziri et al., 2023). In Sec.˜5, we demonstrate that learners may not necessarily generalize on data that are classified as ID in a canonical sense, but coverage can precisely explain when and why this occurs. We also emphasize that coverage is a property of a dataset and is independent of model architectures and learning algorithms, and we demonstrate that the predictions made by our framework are invariant across model architecture and scale in Sec.˜6 and 7. Finally, $k$-coverage can be algorithmically determined for any fixed-length discrete sequence tasks (Alg.˜1), which we use for the analyses in the following sections.
@@ 4 Experimental Setup
Dataset construction.   We construct four synthetic tasks with different structures: 2-Hop, Parallel 2-Hop, 3-Hop, and Non-tree (Fig.˜2). To isolate functional-equivalence-based generalizations, we randomly create ground-truth mappings from a product space of token sets to control the generalization sources not attributable to compositional structures (i.e., commutativity). We explain the dataset construction process using 2-Hop task (Fig.˜2(a)), $(x_{1},x_{2},x_{3})\mapsto t$ with $t=f_{2}(f_{1}(x_{1},x_{2}),x_{3})$, as an example. We construct training datasets by defining a token set with size $\left|{\mathcal{X}}\right|$ and randomly defining two mappings for the primitive functions $f_{1}:{\mathcal{X}}^{2}\to{\mathcal{X}}$ and $f_{2}:{\mathcal{X}}^{2}\to{\mathcal{X}}$. We mark a fraction $p_{\mathrm{seen}}=0.7$ of each function’s domain as ‘seen’, gather all possible combinations where both functions are applied to inputs from their seen domains, and uniformly sample $N$ examples to form a training dataset. See Sec.˜B.1 for more details of the dataset construction process.
Training & evaluation.   Following Wang et al. (2023), we train randomly initialized GPT-2 (Radford et al., 2019) models with 8 layers, 12 heads, and 768 dimensions as a base model (see Sec.˜B.2 for details). We construct two evaluation sets, each with 2,000 instances: (1) ID Test Set : all primitive function applications (e.g., $f_{1}(x_{1},x_{2})$ and $f_{2}(b,x_{3})$ in 2-Hop task) are observed during training, but their specific combination was unseen. (2) Out-of-coverage (canonical OOD) Test Set: at least one primitive function application is never observed during training, which is used as a control group.
@@ 5.1 Evidence Strength is Tightly Aligned to Pattern-Matching Success
We first analyze the correlation between $k$-coverage and ID generalization performance of the GPT-2 model. To this end, we implement and release a task-agnostic coverage determination algorithm (see App.˜C) that can be applied to diverse compositional structures. Then, we analyze what fraction of ID test data of 2-Hop task with $\left|{\mathcal{X}}\right|=50$ lies inside $k$-coverage, depending on $k$ and dataset size $N$. Fig.˜3 (Left) shows that, at N=5N=5k, every ID test example is already covered with minimal evidence ($k=1$). Hence, in an ideal scenario where a single witness of functional equivalence suffices, training with the dataset as small as N=5N=5k will lead to perfect ID generalization.
@@
These results yield two important insights. First, successful ID generalization in practice requires a robust coverage so the model can confidently identify and utilize functional equivalence relationships. The parameter $k$ effectively quantifies this evidence strength, directly impacting generalization speed and reliability. Second, while our experiments use uniformly sampled datasets, the results can explain why models struggle with generalizing long-tail distributions in imbalanced real-world data (Mallen et al., 2023; Kandpal et al., 2023; Chang et al., 2024). Despite technically being in-distribution, rare combinations naturally receive limited evidence of functional equivalence (low $k$), effectively and practically placing them outside the coverage. We believe that our insights will motivate future research on targeted data augmentation strategies to maximize $k$-coverage.
@@ 5.2 Latent Representation Clusters Drive Pattern Matching on Coverage
To quantify this representation clustering phenomenon, we develop a metric that captures how distinctly the model separates functionally equivalent fragments from others. Specifically, we measure the difference between the average pairwise cosine similarity of latent vectors that share the same intermediate state $b=f_{1}(x_{1},x_{2})$ ($\overline{\cos}_{\text{intra}}$), and those that do not ($\overline{\cos}_{\text{inter}}$), for each position and layer of the model. We term this difference the intra–inter cosine gap $\text{IICG}=\overline{\cos}_{\text{intra}}-\overline{\cos}_{\text{inter}}$, where higher values indicate stronger within-group clustering relative to between-group separation. Fig.˜4 (Left) reveals a positive correlation: higher $k$-cutoff values yield higher IICG scores at certain positions, indicating that stronger functional equivalence evidence leads to more coherent internal representations. In contrast, out-of-coverage ($k=0$) examples exhibit no clustering pattern, as they lack evidence of functional equivalence in the training data. The PCA visualization at position $x_{2}$ and layer 3 (Right) shows this trend visually. We verify that the representation clusters play a causal role in pattern matching with causal tracing (Goldowsky-Dill et al., 2023; Hanna et al., 2023), a widely used technique to identify Transformer circuits (Fig.˜8).
Our findings extend the previous insights from mechanistic interpretability literature (Wang et al., 2024a) in several ways. First, we demonstrate that unified circuit formation is driven by functional equivalence evidence in the training data, not by explicit exposure to intermediate computation steps. Moreover, we find that these clustered representations are not necessarily aligned with vocabulary embeddings, implying that standard interpretability methods like logit lens (nostalgebraist, 2020) may fail to detect these functional equivalence representations despite their presence (see App.˜E).